### v2

@@ 2 Setting and main result
[…]
For the operator family $\mathbf{R}=\{R_{\omega}\}_{\omega\in\Omega}$, we need the conditions (8) and (10). We also assume that the operator $P_{\rho}=\mathbb{E}(R_{\omega}R_{\omega}^{\ast})$ is compact. Finally, we assume (14) and that $u\in V_{P_{\rho}}^{s}$ for some $0<s\leq 1$.
Consider the online learning algorithm (4), where $u^{(0)}\in V$ is arbitrary, the parameters $\alpha_{m},\mu_{m}$ are given by (6) with $t=t_{s}:=(1+s)/(2+s)$ and $A=1/(2\Lambda)$, and the random samples $(\omega_{m},y_{m})$, $m=0,1,\ldots,N\leq\infty$, are i.i.d. with respect to $\mu$. Then the expected squared error in $V$ satisfies
@@
In this generality, Theorem 2.1 has not yet appeared in the literature, at least to our knowledge. Its proof is given in Section 4. For the parameter range $0<s\leq 1$, the exponent $-s/(2+s)$ in the right-hand side of (17) is best possible under the general conditions stated in Theorem 2.1. Estimates of the form (17) also hold for arbitrary values $11/2<t<$ 1 and $0<A\leq 1/(2\Lambda)$ in (6), albeit with non-optimal exponents depending on $t$ and different constants $C$ varying with $t$ and $A$. Without the condition (14), which ensures the existence of the minimizer $u\in V$ in (7), the online method (4)-(6) diverges in expectation. Note that convergence estimates for with respect to the weaker $L^{2}_{\rho}(\Omega,Y)$ norm are also of great interest, but cannot be obtained within our framework. We will comment on these issues in the concluding Section 5.
@@
There is a huge amount of literature devoted to the convergence theory of various versions of the algorithm (4)-(5), especially for the scalar-valued case $Y=\mathbb{R}$. In particular, it is often considered in the so-called finite horizon case, where $N<\infty$ is fixed and the step-sizes $\mu_{m}$ are chosen in dependence on $N$ so that expectations such as $\mathbb{E}(\|u-u^{(N)}\|_{V}^{2})$ or $\mathbb{E}(\|f_{u}-f_{u^{(N)}}\|_{L^{2}_{\rho}(\Omega,Y)}^{2})$, respectively, are optimized for the final approximation $u^{(N)}$. A brief discussion of known results is given in the next section.
3 Results related to Theorem 2.1
@@
Given the large number of publications on convergence rates for learning algorithms, we will present only a selection of results focusing on the RKHS setting and online algorithms similar to (4)-(5). The results we cite are often stated and proved for the scalar-valued case $Y=\mathbb{R}$, although some authors claim that their methods extend to the case of an arbitrary separable Hilbert space $Y$ with minor modifications. One of the first papers on the vector-valued case is CaVi2007 , where the authors provide upper bounds in probability for the $L^{2}_{\rho}(\Omega,Y)$ error of $f_{u_{N}}$ if $N\to\infty$ and $\kappa_{N}\to 0$, where $u_{N}$ is the solution of (15). These bounds depend in a specific way on the smoothness of $u\in V_{P_{\rho}}^{s}$, $0\leq s\leq 1$, and on the spectral properties of $P_{\rho}$. Note that in CaVi2007 and in many other papers stronger assumptions on the compactness of $P_{\rho}$ are made than we do, and that bounds in probability do not automatically imply bounds in expectation. Moreover, the error measured in the $L^{2}_{\rho}(\Omega,Y)$ norm is with respect to $f_{u_{N}}$ and not with respect to approximations such as $f_{u^{(m)}}$, $m\leq N$, which are produced by a special algorithm comparable to (5).
@@
In TaYa2014 , the authors provide estimates in probability for an algorithm similar to (4)-(5) for the scalar-valued case $Y=\mathbb{R}$. They cover both, convergence in $L^{2}_{\rho}(\Omega,\mathbb{R})$ and $H$ norms. There, the main additional assumption needed for the application of certain results from martingale theory is that, for some constant $M_{\rho}<\infty$, the random variable $y$ satisfies
@@
for some constant $C=C(M_{\rho},\|u\|_{V_{P_{\rho}}^{s}},m_{0},s,\Lambda,\log(2/\delta))<\infty$. Here $V=H$ is an RKHS of functions $u:\,\Omega\to\mathbb{R}$ generated by some scalar-valued Mercer kernel $K:\,\Omega\times\Omega\to\mathbb{R}$ and $\Lambda=\max_{\omega\in\Omega}K(\omega,\omega)$. The associated maps $R_{\omega}$ are given by $R_{\omega}y=yK(\omega,\cdot)$, $y\in\mathbb{R}$. Consequently, $R_{\omega}^{\ast}u=u(\omega)$, $\omega\in\Omega$, corresponds to function evaluation. So, for $0<s\leq 1$, we get the same rate as in our Theorem 2.1 which was, however, based on the expectation of the squared error in $V=H$ in the more general vector-valued case. What our rather elementary method does not provide is a result for the case $1<s\leq 2$ and for $L^{2}_{\rho}(\Omega,Y)$ convergence. For the latter situation, (TaYa2014, , Theorem C) gives the better estimate
@@
This is almost matching the lower estimates for kernel learning derived in CaVi2007 for classes of instances where the spectrum of $P_{\rho}$ has a prescribed decay of the form $\lambda_{k}\asymp k^{-b}$ for some $b>1$. Note that for Mercer kernels, the operator
[…]
is trace class, while in our Theorem 2.1 no stronger decay of eigenvalues is assumed.
Estimates in expectation close to our result have also been obtained for slightly different settings. For example, in YiPo2008 both, the so-called regularized ($\alpha_{m}<1$) and the unregularized online algorithm ($\alpha_{m}=1$) were analyzed in the scalar-valued case $Y=\mathbb{R}$ under assumptions similar to ours for $L^{2}_{\rho}(\Omega,\mathbb{R})$ and $V=H$ convergence. We quote only the result for convergence in the RKHS $V=H$. It concerns the so-called finite horizon case of the unregularized online algorithm (4) with $\alpha_{m}=1$, where one fixes $N<\infty$, chooses a constant step-size $\mu_{m}=\mu$, $m=0,\ldots,N-1$, which depends on $N$, stops the iteration at $m=N$ and asks for a good estimate of the expectation of $\mathbb{E}(\|u-u^{(N)}\|^{2}_{V})$ for the final iterate only. Up to the specification of constants, Theorem 6 in YiPo2008 states that, under the condition $u\in V^{s}_{P_{\rho}}$, $s>0$, one can obtain the bound
@@
if one sets $\mu_{N}=cN^{-(s+1)/(s+2)}$ with a properly adjusted value of $c$. Note that $s>0$ is arbitrary with the exponent approaching $-1$ if the smoothness parameter $s$ tends to $\infty$, while our result gives no improvement for $s>1$. The drawback of the finite horizon case is that the estimate concerns only a fixed iterate $u^{(N)}$ with an $N$ to be decided beforehand. In a sense, this can be seen as building an approximation to the solution $u_{N}$ of (15) with $\kappa_{N}=\mu_{N}$ from a single pass over the $N$ i.i.d. samples $(\omega_{m},y_{m})$, $m=0,\ldots,N-1$.
@@
on the eigenvalues of the covariance operator $P_{\rho}$. Bounds that require knowledge of $\beta<1$ are sometimes called capacity dependent, so our bounds in Theorem 2.1 as well as the cited results from TaYa2014 ; YiPo2008 are capacity independent. Capacity dependent convergence rates for the expected squared error for the online algorithm (4)-(5) have been obtained, among others, in DiBa2016 ; DFB2017 ; GuSh2019 ; GLZ2022 , again in the scalar-valued case $Y=\mathbb{R}$ and with various parameter settings, including unregularized and finite horizon versions. In DiBa2016 , rates for $\mathbb{E}(\|u-\bar{u}^{(m)}\|_{L^{2}_{\rho}(\Omega,\mathbb{R})}^{2})$ have been established, where
$\bar{u}^{(m)}=\frac{1}{m+1}\sum_{k=0}^{m}u^{(k)},\qquad m=0,1,\ldots,$ (19)
@@
where $u^{(m+1)}=u^{(m)}+\mu_{m}(y_{m}-R_{\omega_{m}}u^{(m)})$, and compares with the regularized iteration (4) with $\alpha_{m}$ given in (6). To illustrate the influence of $\beta$, we formulate the following bound, which is a consequence of (DiBa2016, , Corollary 3): Under an additional technical assumption on the noise term $\epsilon_{\omega}$, if the condition (18) holds for some $10<\beta<$ 1 and $u\in V_{P_{\rho}}^{s}$, $s>-1$, then for suitable choices for the learning rates $\mu_{m}$, we have
$\mathbb{E}(\|u-\bar{u}^{(m)}\|_{L^{2}_{\rho}(\Omega,\mathbb{R})}^{2})=\left\{\begin{array}[]{ll}\mathrm{O}((m+1)^{-(s+1)}),&-1<s<-\beta,\\ \mathrm{O}((m+1)^{-(s+1)/(s+1+\beta)}),&-\beta<s<1-\beta,\\ \mathrm{O}((m+1)^{-(1-\beta/2)}),&1-\beta<s.\end{array}\right.$
Thus, stronger eigenvalue decay generally implies stronger asymptotic error decay in the $L^{2}_{\rho}(\Omega,\mathbb{R})$ norm. In (DFB2017, , Section 6), similar rates are obtained in the finite horizon setting for both, the above averaged iterates $\bar{u}^{(N)}$ and for $u^{(N)}$ produced by a two-step extension of the one-step iteration (4).
In addition to $L^{2}_{\rho}(\Omega,\mathbb{R})$ convergence results, the paper GuSh2019 also provides a capacity dependent convergence estimate in the RKHS norm for the unregularized algorithm (4)-(5) with parameters $\alpha_{m}=1$ and $\mu_{m}=c(m+1)^{-1/2}$. Under the boundedness assumption $|y|\leq M_{\rho}$, Theorem 2 in GuSh2019 implies that
$\mathbb{E}(\|u-{u}^{(m)}\|^{2}_{V})=\mathrm{O}((m+1)^{-\min(s,1-\beta)/2)}\log^{2}(m+1)),\qquad m=1,2,\ldots,$
if $u\in V^{s}_{P_{\rho}}$ for some $s>0$, $P_{\rho}^{\beta}$ is trace class for some $10<\beta<$ 1, and $c$ is properly adjusted.
Finally, the scalar-valued least-squares regression problem with $Y=\mathbb{R}$ and RKHS prior space $H$ can also be cast as linear regression problem in $V=H$. This was done in DFB2017 ; GLZ2022 . More abstractly, given a $\mu$-distributed random variable $(\xi_{\omega},y)\in V\times\mathbb{R}$ on $\Omega\times\mathbb{R}$, the task is to find approximations to the minimizer $u\in V$ of the problem
$\mathbb{E}(|(\xi_{\omega},v)-y|^{2})\longmapsto\min,\qquad v\in V,$ (20)
from i.i.d. samples $(\xi_{\omega_{i}},y_{i})$. If $V=H$ is the RKHS that regularizes the scalar-valued least-squares regression problem on $\Omega\times\mathbb{R}$, then the canonical choice is $\xi_{\omega}=K(\omega,\cdot)$. In GLZ2022 , for the iteration
@@
weak convergence in $V$ is studied by deriving estimates for quantities such as $\mathbb{E}((v,u-u^{(m)})^{2})$ and $\mathbb{E}((\xi_{\omega^{\prime}},u-u^{(m)})^{2})$ under some simplifying assumptions on the noise and the normalization $\|\xi_{\omega}\|=1$. Note that this iteration is nothing but the unregularized iteration (4) with $\alpha_{m}=1$ since $(\xi_{\omega_{m}},u^{(m)})=u^{(m)}(\omega_{m})$ in this case. In the learning application, the assumption $\|\xi_{\omega}\|=1$ means $K(\omega,\omega)=1$. Moreover, in this case
$\mathbb{E}((\xi_{\omega^{\prime}},u-u^{(m)})^{2})=\mathbb{E}(\|u-u^{(m)}\|_{L^{2}_{\rho}(\Omega,\mathbb{R})}),$
since the expectation on the left, in addition to the i.i.d. samples $(\xi_{\omega_{k}},y_{k})$, $k=0,\ldots,m-1$, is also taken with respect to the independently $\rho$-distributed random variable $\xi_{\omega^{\prime}}$. This is related to learning rates in the $L^{2}_{\rho}(\Omega,\mathbb{R})$ norm. The estimates for $\mathbb{E}((\xi_{\omega^{\prime}},u-u^{(m)})^{2})$ given in GLZ2022 concern both, the finite horizon and the online setting and again depend on the parameters $s\geq 0$ (smoothness of $u$) and $0<\beta\leq 1$ (capacity assumption on $P_{\rho}$). For the estimates of $\mathbb{E}((v,u-u^{(m)})^{2})$, the smoothness $s^{\prime}\geq 0$ of the fixed element $v\in V^{s^{\prime}}_{P_{\rho}}$ is traded against the smoothness $s\geq 0$ of $u\in V^{s}_{P_{\rho}}$. See GLZ2022 for details.
@@
[…]
@@
under the assumption $u\in V_{P_{\rho}}^{s}$, $0<s\leq 1$, if the parameters $A$ and $t$ in (6) are chosen accordingly. The precise statement and the dependence of the constant in (21) on the initial error, the noise variance and the smoothness assumption are given in the formulation of Theorem 2.1.
@@
where $\bar{\alpha}_{m}:=1-\alpha_{m}=(m+2)^{-1}$, compare also 6. The first term $\bar{e}^{(m+1)}$ corresponds to the noiseless case considered in GrOs2018 , while the remainder term is the noise contribution. Thus,
@@
Here the term involving $(e^{(m)},P_{\rho}e^{(m)})\geq 0$ has been omitted, since its forefactor $-\mu_{m}(1-2\Lambda\mu_{m})$ is non-positive due to the restriction on $A$ in (6).
@@
Next, we switch to full expectations in (27) by using the independence assumption for the sampling process and taking into account that
@@
Since $11/2<t<$ 1 is assumed in (6), we have $a>0$. Then, for $k=0,1,\ldots,m$, the estimate for $\varepsilon_{k+1}$ simplifies to

### v5

@@ 2 Setting and main result
$\mathbb{E}(\|y\|_{Y}^{2})=\int_{\Omega\times Y}\|y\|_{Y}^{2}\,d\mu<\infty.$
For the feature map $\mathbf{R}=\{R_{\omega}\}_{\omega\in\Omega}$, we require uniform boundedness (7) and measurability. We also assume that the operator $P_{\rho}=\mathbb{E}(R_{\omega}R_{\omega}^{\ast})$ is injective and compact. Finally, we assume (10) and that $u\in V_{P_{\rho}}^{s}$ for some $0<s\leq 1$.
Consider the online learning algorithm (3), where $u^{(0)}\in V$ is arbitrary, the parameters $\alpha_{m},\mu_{m}$ are given by (5) with $t=t_{s}:=(1+s)/(2+s)$ and $A=1/(2\Lambda)$, and the random samples $(\omega_{m},y_{m})$, $m=0,1,\ldots,N\leq\infty$, are i.i.d. with respect to $\mu$. Then the expected squared error satisfies
@@
In this generality, Theorem 2.1 has not yet appeared in the literature, at least to our knowledge. Its proof is given in Section 4. For the parameter range $0<s\leq 1$, the exponent s/(2+s)-s/(2+s) in the right-hand side of (13) is best possible under the general conditions stated in Theorem 2.1, compare CaVi2007 ; LGZ2017 . Estimates of the form (13) also hold for arbitrary values $1/2<t<1$ and $0<A\leq 1/(2\Lambda)$ in (5), albeit with non-optimal exponents depending on $t$ and different constants $C$ varying with $t$ and $A$. Without the condition (10), which ensures the existence of the minimizer $u\in V$ in (6), the online method (3)-(5) diverges in expectation. Note that convergence estimates with respect to the weaker $L^{2}_{\rho}(\Omega,Y)$ norm (hard learning scenario) cannot be obtained within our framework. We will comment on these issues in the concluding Section 5.
@@
There is a huge amount of literature devoted to the convergence theory of various versions of the algorithm (3)-(4), especially for the scalar-valued case $Y=\mathbb{R}$. In particular, the algorithm is often considered in the so-called finite horizon case, where $N<\infty$ is fixed and the step-sizes $\mu_{m}$ are chosen in dependence on $N$ so that expectations such as $\mathbb{E}(\|u-u^{(N)}\|_{V}^{2})$ or $\mathbb{E}(\|f_{u}-f_{u^{(N)}}\|_{L^{2}_{\rho}(\Omega,Y)}^{2})$, respectively, are optimized for the final approximation $u^{(N)}$. A brief discussion of known results is given in the next section.
3 Examples and results related to Theorem 2.1
First, we provide some examples for the framework of this paper. We start with the scalar-valued case $Y=\mathbb{R}$ (or $Y=\mathbb{C}$), where suitable separable RKHS $H$ of functions $f:\,\Omega\to\mathbb{R}$ are often directly given by their associated bounded, symmetric, positive definite kernel function $k:\,\Omega\times\Omega\to\mathbb{R}$. Concrete examples of kernels, including Sobolev and Gaussian kernels, can be found in StCh2008 ; PaRa2016 . The canonical choice for the feature space $V$ is to identify $V=H$ and to define the feature map by
$R_{\omega}y=k(\omega,\cdot)y,\qquad R_{\omega}^{\ast}v=(k(\omega,\cdot),v)_{H}=v(\omega),\qquad y\in\mathbb{R},\quad v\in H,\quad\omega\in\Omega.$
As expected, the operator $P_{\rho}$ coincides with the integral operator
$(P_{\rho}v)(\omega)=\int_{\Omega}k(\omega,\theta)v(\theta)\,d\rho(\theta).$
We just note that the assumption (10) in Theorem 2.1 is equivalent to
$\int_{\Omega}k(\omega,\theta)f_{\mu}(\theta)\,d\rho(\theta)\in\mathrm{ran}(P_{\rho}),$
which essentially means that $f_{\mu}$ must coincide with an element $v$ in the RKHS $H$ up to an additive perturbation from $\mathrm{ker}(\tilde{P}_{\rho})$. Here $\tilde{P}_{\rho}$ denotes the extension of $P_{\rho}$ to all of $L^{2}_{\rho}(\Omega,\mathbb{R})$. In the literature, authors usually assume $f_{\mu}\in H=V_{P_{\rho}}^{0}$. The choices of feature space and map are by no means unique. For example, one can set $V=\ell^{2}(\mathbb{N})$, fix an arbitrary complete orthonormal system (CONS) $(\phi_{i})_{i\in\mathbb{N}}$ in $H$ and define the feature map by
$R_{\omega}y=(y\phi_{i}(\omega))_{i\in\mathbb{N}},\qquad R_{\omega}^{\ast}c=\sum_{i\in\mathbb{N}}c_{i}\phi_{i}(\omega),\qquad y\in\mathbb{R},\quad c:=(c_{i})_{i\in\mathbb{N}}\in\ell^{2}(\mathbb{N}),\quad\omega\in\Omega.$
We do not go into further detail. Another example with $Y=\mathbb{R}$, where the RKHS $H$ can be identified with $\ell^{2}(\mathbb{N})$ will be considered in Subsection 5.4.
The construction of operator kernels $K$ for vector-valued RKHS typically involves one or several scalar-valued kernels $k$. Here, we present an example of a multiplicative (or separable) operator kernel
$K(\omega,\theta)=k(\omega,\theta)T,$ (14)
where $T:\,Y\to Y$ is a bounded, positive definite, selfadjoint operator, which is a modification of (KDPCRA2015, , Example 3). Let $Y=H^{1}_{0}(G)$, where $G$ is a fixed domain in $\mathbb{R}^{d}$ and set for simplicity $\Omega=[0,1]$. The associated learning problem is about finding a regression function $f_{\mu}:\,[0,1]\to H_{0}^{1}(G)$ with values in an infinite-dimensional Sobolev space. [^7] A potential application could be finding approximations to the solution map $\omega\to u_{\omega}$ for the diffusion problem $\nabla(a_{\omega}\nabla u)=f$ with random diffusion coefficient $a_{\omega}$, fixed right-hand side $f$ and solution $u=u_{\omega}\in H^{1}_{0}(G)$ from samples $(\omega_{i},u_{i}\approx u_{\omega_{i}})$. Note that $Y$ could be any separable Hilbert space in the considerations below. For the feature space, consider the choice $V=L^{2}(\Omega)\otimes Y$ (this space can be identified with the Bochner space $L^{2}(\Omega,Y)$) and introduce the feature map by
$R_{\omega}u=(\chi_{[0,\omega)}-\omega)\otimes u,\qquad\omega\in\Omega,\quad u\in Y,$
where $\chi_{E}$ denotes the indicator function of a set $E\subset\Omega$. Then, on elementary tensors, the adjoints $R_{\omega}^{\ast}$ are given by
$R_{\omega}^{\ast}(f\otimes v)=(\int_{0}^{\omega}f(t)\,dt-\omega\int_{0}^{1}f(t)\,dt)v,\qquad f\in L^{2}(\Omega),\quad v\in Y.$
It takes some basic calculations to check that the associated operator kernel is of the form
$K(\omega,\theta)=k(\omega,\theta)I,$
where $I$ is the identity on $Y$ and
$k(\omega,\theta)=\left\{\begin{array}[]{ll}(1-\omega)\theta,&\;\theta\leq\omega,\\ (1-\theta)\omega,&\;\theta\geq\omega,\end{array}\right$.
is the kernel for the scalar-valued RKHS $H_{0}^{1}(\Omega)$. Moreover, the resulting RKHS $H$ can be identified with $H^{1}_{0}(\Omega)\otimes Y$. Roughly speaking, functions in $H$ are of the form
$f(\omega)=\sum_{i\in\mathbb{N}}c_{i}(\omega)\phi_{i}\quad\mbox{with }c_{i}\in H^{1}_{0}(\Omega),$
while for functions in $L^{2}(\Omega,Y)$ such representations require only $c_{i}\in L_{2}(\Omega)$. Here, $(\phi_{i})_{i\in\mathbb{N}}$ is a CONS in $Y$. In other words, convergence in $H$ implies a certain control of the smoothness of the coefficients $c_{i}(\omega)$ as functions on $\Omega$.
@@
Given the large number of publications on convergence rates for learning algorithms, we will present only a selection of results focusing on the RKHS setting and online algorithms similar to (3)-(4). The results we cite are often stated and proved for the scalar-valued case $Y=\mathbb{R}$, although some authors claim that their methods extend to the case of an arbitrary separable Hilbert space $Y$ with minor modifications. One of the first papers on the vector-valued case is CaVi2007 , where the authors provide upper bounds in probability for the $L^{2}_{\rho}(\Omega,Y)$ norm of $f_{u}-f_{u_{N}}$, if $N\to\infty$ and $\kappa_{N}\to 0$, where $u_{N}$ is the solution of (11). These bounds depend in a specific way on the smoothness of $u\in V_{P_{\rho}}^{s}$, $0\leq s\leq 1$, on compactness assumptions for the feature map $R_{\omega}$ and on the spectral properties of $P_{\rho}$. Note that the error measured in the $L^{2}_{\rho}(\Omega,Y)$ norm is with respect to $f_{u_{N}}$ and not with respect to approximations such as $f_{u^{(m)}}$, $m\leq N$, which are produced by a special algorithm comparable to (4). The results in CaVi2007 have recently been extended in FiSt2020 ; LMMG2024 to RKHS with multiplicative kernels (14) to cover $V_{P_{\rho}}^{t}$ error bounds in probability under the assumption $f_{\mu}\in V_{P_{\rho}}^{s}$, $1\leq t<s\leq 1$. This covers $H$ ($t=0$) and $L_{\rho}^{2}(\Omega,Y)$ (t=1t=-1) convergence, including the hard learning scenario $f_{\mu}\not\in H$. The bounds require Bernstein-type inequalities for the noise level which hold in particular for uniformly bounded noise.
@@
In TaYa2014 , the authors provide estimates in probability for an algorithm similar to (3)-(4) in the scalar-valued case $Y=\mathbb{R}$. They cover both, convergence in $L^{2}_{\rho}(\Omega,\mathbb{R})$ and $H$ norms. There, the main additional assumption needed for the application of certain results from martingale theory is that, for some constant $M_{\rho}<\infty$, the random variable $y$ satisfies
@@
for some constant $C=C(M_{\rho},\|u\|_{V_{P_{\rho}}^{s}},m_{0},s,\Lambda,\log(2/\delta))<\infty$. Here $V=H$ is an RKHS of functions $u:\,\Omega\to\mathbb{R}$ generated by some continuous scalar-valued Mercer kernel $k:\,\Omega\times\Omega\to\mathbb{R}$ and $\Lambda=\max_{\omega\in\Omega}k(\omega,\omega)$. So, for $0<s\leq 1$, we get the same rate as in our Theorem 2.1 which is, however, concerned with the expectation of the squared RKHS error in the more general vector-valued case. What our rather elementary method does not provide is a result for the case $1<s\leq 2$ and for $L^{2}_{\rho}(\Omega,Y)$ convergence. For the latter situation, (TaYa2014, , Theorem C) gives the better estimate
@@
This is almost matching the lower estimates for kernel learning derived in CaVi2007 for classes of instances where the spectrum of $P_{\rho}$ has a prescribed decay of the form $\lambda_{k}\asymp k^{-b}$ for some $b>1$, see also LGZ2017 . Recall that the integral operator $P_{\rho}$ is trace class for scalar-valued Mercer kernels, while in our Theorem 2.1 no stronger decay of eigenvalues is assumed. The above cited bounds in probability automatically imply similar bounds for the expectation of squared errors.
Estimates in expectation close to our result have also been obtained for slightly different settings. For example, in YiPo2008 both, the so-called regularized ($\alpha_{m}<1$) and the unregularized online algorithm ($\alpha_{m}=1$) were analyzed in the scalar-valued case $Y=\mathbb{R}$ under assumptions similar to ours regarding $L^{2}_{\rho}(\Omega,\mathbb{R})$ and $V=H$ convergence. We quote only the result for convergence in the RKHS $V=H$. It concerns the so-called finite horizon case of the unregularized online algorithm (3) with $\alpha_{m}=1$, where one fixes $N<\infty$, chooses a constant step-size $\mu_{m}=\mu_{N}$, $m=0,\ldots,N-1$, which depends on $N$, stops the iteration at $m=N$, and asks for a good estimate of the expectation of $\mathbb{E}(\|u-u^{(N)}\|^{2}_{V})$ for the final iterate only. Up to the specification of constants, Theorem 6 in YiPo2008 states that, under the condition $u\in V^{s}_{P_{\rho}}$, $s>0$, one can obtain the bound
@@
if one sets $\mu_{N}=cN^{-(s+1)/(s+2)}$ with a properly adjusted value of $c$. Note that $s>0$ is arbitrary with the exponent approaching 1-1 if the smoothness parameter $s$ tends to $\infty$, while our result provides no improvement for $s>1$. The drawback of the finite horizon case is that the estimate only concerns a fixed iterate $u^{(N)}$ with an $N$ to be decided beforehand. In a sense, this can be seen as building an approximation of the solution $u_{N}$ of (11) with $\kappa_{N}=\mu_{N}$ from a single pass over the $N$ i.i.d. samples $(\omega_{m},y_{m})$, $m=0,\ldots,N-1$.
@@
on the eigenvalues of the operator $P_{\rho}$. Bounds that incorporate knowledge of $\beta<1$ are sometimes called capacity dependent, so our bound in Theorem 2.1 as well as the cited results from TaYa2014 ; YiPo2008 are capacity independent (in contrast, CaVi2007 ; FiSt2020 ; LMMG2024 all deal with capacity dependent estimates). Capacity dependent convergence rates for the expected squared error for the online algorithm (3)-(4) have been obtained, among others, in DiBa2016 ; DFB2017 ; GuSh2019 ; GLZ2022 , again in the scalar-valued case $Y=\mathbb{R}$ ($V=H$) and with various parameter settings, including unregularized and finite horizon versions. In DiBa2016 , rates for $\mathbb{E}(\|u-\bar{u}^{(m)}\|_{L^{2}_{\rho}(\Omega,\mathbb{R})}^{2})$ have been established, where
$\bar{u}^{(m)}=\frac{1}{m+1}\sum_{k=0}^{m}u^{(k)},\qquad m=0,1,\ldots,$
@@
where $u^{(m+1)}=u^{(m)}+\mu_{m}(y_{m}-R_{\omega_{m}}u^{(m)})$, and compares with the regularized iteration (3) with $\alpha_{m}$ given in (5). To illustrate the influence of $\beta$, we formulate the following bound, which is a consequence of (DiBa2016, , Corollary 3): Under an additional technical assumption on the noise term $\epsilon_{\omega}$, if the condition (15) holds for some $0<\beta<1$ and $u\in V_{P_{\rho}}^{s}$, s>1s>-1, then for suitable choices for the learning rates $\mu_{m}$, we have
$\mathbb{E}(\|u-\bar{u}^{(m)}\|_{L^{2}_{\rho}(\Omega,\mathbb{R})}^{2})=\left\{\begin{array}[]{ll}\mathrm{O}((m+1)^{-(s+1)}),\quad&-1<s<-\beta,\\ \mathrm{O}((m+1)^{-(s+1)/(s+1+\beta)}),\quad&-\beta<s<1-\beta,\\ \mathrm{O}((m+1)^{-(1-\beta/2)}),\quad&1-\beta<s.\end{array}\right$.
Thus, stronger eigenvalue decay implies stronger asymptotic error decay in the $L^{2}_{\rho}(\Omega,\mathbb{R})$ norm. In (DFB2017, , Section 6), similar rates are obtained in the finite horizon setting for both, the above averaged iterates $\bar{u}^{(N)}$ and for $u^{(N)}$ produced by a two-step extension of the one-step iteration (3).
In addition to $L^{2}_{\rho}(\Omega,\mathbb{R})$ convergence results, GuSh2019 also provides a capacity dependent convergence estimate in the RKHS norm for the unregularized algorithm (3)-(4) with parameters $\alpha_{m}=1$ and $\mu_{m}=c(m+1)^{-1/2}$. Under the boundedness assumption $|y|\leq M_{\rho}$, Theorem 2 in GuSh2019 implies that
$\mathbb{E}(\|u-{u}^{(m)}\|^{2}_{V})=\mathrm{O}((m+1)^{-\min(s,1-\beta)/2}\log^{2}(m+1)),\qquad m=1,2,\ldots,$
if $u\in V^{s}_{P_{\rho}}$ for some $s>0$, $P_{\rho}^{\beta}$ is trace class for some $0<\beta<1$, and $c$ is properly adjusted.
Finally, the scalar-valued kernel regression problem with $Y=\mathbb{R}$ and prior RKHS $V=H$ can also be cast as linear regression problem in $V=H$. This was done in DFB2017 ; GLZ2022 . More abstractly, given a $\mu$-distributed random variable $(\xi_{\omega},y)\in V\times\mathbb{R}$ on $\Omega\times\mathbb{R}$, the task is to find approximations to the minimizer $u\in V$ of the problem
$\mathbb{E}(|(\xi_{\omega},v)_{V}-y|^{2})\longmapsto\min,\qquad v\in V,$
from i.i.d. samples $(\xi_{\omega_{i}},y_{i})$. If $k$ is the scalar-valued kernel of the RKHS $V=H$ then the canonical choice is $\xi_{\omega}=k(\omega,\cdot)$. In GLZ2022 , weak convergence in $V$ is studied for the iteration
@@
by deriving estimates for quantities such as $\mathbb{E}((v,u-u^{(m)})_{V}^{2})$ and $\mathbb{E}((\xi_{\omega^{\prime}},u-u^{(m)})_{V}^{2})$ under some assumptions on the learning rates $\mu_{m}$, a more restrictive noise model, and the normalization $\|\xi_{\omega}\|_{V}=1$. This iteration is nothing but the unregularized iteration (3) with $\alpha_{m}=1$, since $(\xi_{\omega_{m}},u^{(m)})_{V}=u^{(m)}(\omega_{m})$ in the scalar-valued case. Note that the assumption $\|\xi_{\omega}\|_{V}=1$ means $k(\omega,\omega)=1$. Moreover, in this case
$\mathbb{E}((\xi_{\omega^{\prime}},u-u^{(m)})_{V}^{2})=\mathbb{E}(\|u-u^{(m)}\|_{L^{2}_{\rho}(\Omega,\mathbb{R})}),$
since the expectation on the left, in addition to the i.i.d. samples $(\xi_{\omega_{i}},y_{i})$, $i=0,\ldots,m-1$, is also taken with respect to the independently $\rho$-distributed random variable $\xi_{\omega^{\prime}}$. This implies learning rates in the $L^{2}_{\rho}(\Omega,\mathbb{R})$ norm. The estimates for $\mathbb{E}((\xi_{\omega^{\prime}},u-u^{(m)})_{V}^{2})$ given in GLZ2022 concern both, the finite horizon and the online setting and again depend on the parameters $s\geq 0$ (smoothness of $u$) and $0<\beta\leq 1$ (capacity assumption on $P_{\rho}$). Convergence in the RKHS norm is studied in the finite horizon case, see (GLZ2022, , Corollary 2.8). For the estimates of $\mathbb{E}((v,u-u^{(m)})_{V}^{2})$, the smoothness $s^{\prime}\geq 0$ of the fixed element $v\in V^{s^{\prime}}_{P_{\rho}}$ is traded against the smoothness $s\geq 0$ of $u\in V^{s}_{P_{\rho}}$, see GLZ2022 for details.
@@
In this subsection we will use the notation and assumptions outlined above, with the only change that the scalar product in $V$ is simply denoted by $(\cdot,\cdot)$ and the associated norm $\|\cdot\|_{V}$ is accordingly denoted by $\|\cdot\|$. We also set $e^{(m)}:=u-u^{(m)}$. We will prove an estimate of the form
@@
under the assumption $u\in V_{P_{\rho}}^{s}$, $0<s\leq 1$, if the parameters $A$ and $t$ in (5) are chosen properly. The precise statement and the dependence of the constant in (16) on the initial error, the noise variance and the smoothness assumption are given in the formulation of Theorem 2.1.
@@
where $\bar{\alpha}_{m}:=1-\alpha_{m}=(m+2)^{-1}$, compare also (5). The first term $\bar{e}^{(m+1)}$ corresponds to the noiseless case considered in GrOs2018 , while the remainder term is the noise contribution. Thus,
@@
Here the term involving $(e^{(m)},P_{\rho}e^{(m)})\geq 0$ has been omitted, since its resulting forefactor $\mu_{m}(1-2\Lambda\mu_{m})$ is non-positive due to the restriction on $A$ in (5).
@@
Next, we switch to full expectations in (22) by using the independence assumption for the sampling process and by taking into account that
@@
Since $1/2<t<1$ is assumed in (5), we have $a>0$. Then, for $k=0,1,\ldots,m$, the estimate for $\varepsilon_{k+1}$ simplifies to